OpenStudy (anonymous):

3(x – 6) = 6 solve for x

OpenStudy (hoblos):

first distribute 3x-18 = 6 add 18 for both sides 3x = 24 divide both sides by 3 x= 8

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

solve: 3 (x - 6) = 6 for x Results: x =8

OpenStudy (callisto):

perhaps x(x-6) = 6 x-6 = 2 x=8 is a faster way

OpenStudy (anonymous):

Thank you guys

OpenStudy (anonymous):

welcome.